Your Pizza is a small pizzeria at Tuskegee, AL. In the table following, the difference in cost structure is summarized. Note that both plants are assumed to have the same labor costs of $400 per week, the same fixed cost of $1,000 per oven per week, and the same opportunity cost of $500 per week. Note that the only difference in conditions facing the two plants is the breakdown of total costs into its fixed and variable components. The firm’s long-run decisions of individual plant depend on what their costs are likely to be at different scales of operation. You must decide to operate Plant 1, Plant 2, or both.

Labor
(workers per week)

Output with Plant 1 (1 oven)
(pizzas per week)

Output with Plant 2 (2 ovens)
(pizzas per week)

1

6

10

2

11

22

3

15

35

4

18

45

5

21

52

6

23

55


Suppose Your Pizza uses Plant 1 and hires 2 workers. What is the firm's total cost and marginal cost?

$2,300 and $400

$1,900 and $800

$3,900 and $1,200

$2,700 and $800

$3,500 and $400
Answer #1

Answer : The answer is option A.

Cost of 1 labor = $400

So, the cost of 2 labors = 400 * 2 = $800

Total cost for plant 1 when hires 2 labors = Total labor cost + Fixed cost + Opportunity cost

=> Total cost for plant 1 when hires 2 labors = 800 + 1,000 + 500 = $2,300

Marginal cost comes from only variable cost. Here the labor cost is the only variable cost. Marginal cost shows the cost of hiring one additional unit of input in production process. As here labor is the only variable input hence hiring one additional labor creates the cost of $400. So, here the marginal cost is $400.

Therefore, option A is correct.